A random sample of 14 cigarettes of a certain brand has an average nicotine content of 3.9 milligrams and a standard deviation of 2.5 milligrams. What is the lower bound of the 99% confidence interval for the average nicotine content of the cigarettes.



1
Expert's answer
2022-02-25T11:06:13-0500

The critical value for "\alpha = 0.01" and "df = n-1 = 13" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =3.012275."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(3.9-3.012275\times\dfrac{2.5}{\sqrt{14}}, 3.9+3.012275\times\dfrac{2.5}{\sqrt{14}})"

"=(1.887, 5.913)"

The lower bound of the 99% confidence interval for the average nicotine content of the cigarettes is 1.887 milligrams.